Calculate the molar solubility of silver chromate in pure water. Ksp = 1.12 x 10¯12
Answer #1

Solubility equilibrium of Ag2CrO4(s) is

Ag2CrO4(s) <--------> 2Ag+(aq) + CrO42-(aq)

Ksp = [Ag+]2[CrO42-] = 1.12 ×10-12

If solubility of Ag2CrO4 is represented by S ,

at saturated solution

[Ag+] = 2S

[CrO42-] = S

(2S)2S = 1.12 ×10-12

4S3 = 1.12 ×10-12

S3 = 2.8 ×10-13

S = 6.54 ×10-5

Therefore

​​​​​​Molar solubility of Ag2CrO4 = 6.54 ×10-5 mol/L